OpenStudy (anonymous):

A particulat style of pants costs the retailer $70 per pair. At what price should the retailer mark them so he can sell them at a 10% discount off the selling price and still make a 35% profit on his cost?

OpenStudy (blacksteel):

Alright, to do this one we need to figure out what price the retailer needs to sell at to make a 35% profit, and then mark that price up by 10%. To make a 35% profit, the retailer must sell them at 1.35*$70, or $94.50. Then we must mark them up 10% so that after applying a 10% discount we're still making that much. Thus our final selling price is 1.1*$94.50 = $103.95.

OpenStudy (anonymous):

Final selling price is $ 103.95??

OpenStudy (anonymous):

$70 from begin or after discount?

OpenStudy (anonymous):

blacksteel has it. the question is ". At what price should the retailer mark them " and the answer is what blacksteel wrote. then the retailer can mark them down 10% and still make a 35% profit.

OpenStudy (blacksteel):

I'm afraid Nancy's answer is incorrect. 103.95 is a 48.5% markup from 70 - you don't need to mark the pants up by 35%, then 10%, then 10% again; you only need to do the 10% markup once (because that's the markup you're removing to meet your 35% profit.)

OpenStudy (anonymous):

start $70 want 35% profit need sell at price $94.5 and mark Price $115 ,than 10% discount of $115 is $94.5

OpenStudy (anonymous):

70+(70*0.35)= 70+ 24.5 = 94.5 need sell $94.4 will make 35% !0% discount x-(x*0.1)= $94.5 0.9x= $94.5 x = $94.5/0.9 x = $105 Check: 105-(105*0.1)=105-10.5 =$ 94.5 Check: 103.95-(103.95*0.1)=103.95-10.395 =$ 93.555 need $94.5 made 35% profit not $93.555 than the correct answer is $105 not $ 103.95